Shift report 20010921 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Sep 21 2001 - 01:14:35 EDT

  • Next message: Apache: "Shift report 20010921 00:00-08:00"

    ------------------------------------------------------------
    Date: Fri Sep 21 01:14:35 2001, Shift: 16:00-24:00
    Supervisor: yklee, with: 
    Log book#9 pages 17 - 19
    
    Shift Summary:
    ------------------------------------------------------------
    1. RHIC started filling at 2200. Since then we had
       good beam.
    2. Run4831 with trigger1=145K, with T3 Anode1 tripping.
       Run4831 with trigger1=80K, with troubled T3,T4,T5.
    	Magnet setting MRS 4A, FS 1/5 A
    3. HV mainframe 4 and 5 has been bypassed by remote control.
       HV default settings for these mainframes were checked and
       control was set to local control during an access to IR
       as an emergency measure. The mainframe 4 and 5 should
       not be recycled until the next access.
       T3 Anode-1 and Cathode-1 could not be turned on. Ramiro
       suspects there is a short in Anode-1 and Cathode-1 of T3.
    4. Run 4831 has pulser Trigger 7, scaled down by 70, with
       with drift velocity monitor pulser so that DVM is now
       integrated into data runs.
    
    Remember that Mainframe 4 and 5 should not be cycled until
    the next access.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Fri Sep 21 2001 - 01:15:15 EDT